# NVILA: Efficient Frontier Visual Language Models

### v1

@@ 3.3 Efficiency Results
We compare NVILA’s inference performance against Qwen2-VL [3] as shown in Figure 6. For a fair comparison, both models process video inputs by sampling 64 frames, with all experiments conducted on a single NVIDIA RTX 4090 GPU. Qwen2-VL is quantized to W4A16 and deployed with vLLM [48], a LLM/VLM serving engine with state-of-the-art inference speed. For NVILA, we quantize the LLM backbone to W4A16 and vision tower to W8A8. With our specialized inference engine, NVILA achieves up to 2.2$\times$ speedup in prefilling stage and up to 2.8$\times$ higher decoding throughput over Qwen2-VL.
@@ 4.1 Temporal Localization
Table 10: Temporal localization on ActivityNet.
Mean IoU Precision @ 0.5
LITA-7B 24.1 21.1
LITA-13B 28.6 25.9
NVILA-8B (Ours) 34.8 32.1
Figure 7: NVILA deployed as a Vision-Language Navigation agent, navigating environments using language instructions and visual observations (Top: simulation, Bottom: real-world). The real-world setup features a Unitree Go2 robot equipped with a LiDAR sensor at the base of its head and an Intel RealSense Camera mounted on top. On the server side, an RTX 4090 GPU powers the NVILA-8B model, configured with an 8-frame context length for action generation.
@@ 4.2 Robotic Navigation
Table 11: Robotic navigation on the R2R Val-Unseen split. We report standard metrics in VLN-CE: Navigation Error (NE), Oracle Success Rate (OS), Success Rate (SR), and Success-weighted Path Length (SPL).
@@
Method NE $\downarrow$ OS $\uparrow$ SR $\uparrow$ SPL $\uparrow$
Seq2Seq-RGB [49] 11.2 12.2 0.0 0.0
CMA-RGB [49] 9.55 14.8 0.0 0.0
NaVid [7] 5.47 49.0 37.0 35.0
NVILA (Ours) 6.72 55.3 42.5 36.6
4.3 Medical Multi-Modal VILA-M3
NVILA also offers transformative potential in the medical domain. Such integration promises advancements in diagnostic accuracy, clinical decision-making, and data interpretation.
The NVILA-M3 framework [9] introduces a novel approach by integrating multiple domain-expert models tailored to specific medical tasks, such as image segmentation and classification (Fig. 8). These expert models are designed to extract and interpret intricate features that are otherwise difficult for general VLM’s to discern. By coupling these specialized models with a vision-language learning paradigm, NVILA-M3 achieves enhanced performance, facilitating the learning of nuanced relationships between visual inputs and their textual annotations. This integration not only improves task-specific outcomes but also sets a foundation for the development of more robust and context-aware VLMs in the healthcare domain. NVILA-M3 indicated that an overall improvement of  9% can be achieved via usage of expert models over existing SOTA, a few key results can be observed in Table. 12. This underscores the importance of leveraging domain expertise to bridge the gap between generalized AI capabilities and the demands of specialized applications, demonstrating the potential for VLMs to revolutionize fields where precision and specificity are paramount.
Figure 8: Overview of M3 depicting integration of domain expert model with NVILA. Expert model cards are information for NVILA to decide which expert model to call and then utilize results before providing a response to the user. Top: Reference model cards can be seen and multiple use-case queries that a user could present such as segmentation and identification of tumor. Bottom: With guidance of expert models M3 can provide more accurate responses
Table 12: Performance of best M3 model on key benchmarks is shown. Task-specific SOTA baselines and datasets are described in the experiments section [9]. Datasets where * is shown for Med-Gemini are either if a subset is used or pre-processing details have not been disclosed. Metrics for VQA is accuracy, for report generation ROUGE score and for classification F1 score have been utilized
Model VQA Rep. Gen. Classif.
Dataset Rad Path CXR CheXpert
Med-Gemini 78.8* 83.3 28.3* 48.3
Task Spcf. SOTA 84.2 91.7 30.6 51.5
NVILA-M3 (ours) 90.4 92.7 32.2 61.0
@@ 5.1 Visual Language Models
Meanwhile, open-source VLMs continue to evolve, improving at both the system/framework level [59] and the algorithm/recipe level [2], progressively narrowing the performance gap between proprietary and open-source models [17, 60, 61, 62]. These recent advancements have led many open VLM models to claim performance levels comparable to, or even exceeding, leading proprietary models such as GPT-4V and GPT-4o. Some representative examples include InternVL2 [4], Qwen2-VL [3], LLaVA-OneVision [5], Llama 3.2 Vision [63], Molmo [64], NVLM [62], and MiniCPM-V [16].

### v2

@@ 3.3 Efficiency Results
We compare NVILA’s inference performance against Qwen2-VL [5] as shown in Figure 5. For a fair comparison, both models process video inputs by sampling 64 frames, with all experiments conducted on a single NVIDIA RTX 4090 GPU. Qwen2-VL is quantized to W4A16 and deployed with vLLM [64], a LLM/VLM serving engine with state-of-the-art inference speed. For NVILA, we quantize the LLM backbone to W4A16 and vision tower to W8A8. With our specialized inference engine, NVILA achieves up to 2.2$\times$ speedup in pre-filling stage and up to 2.8$\times$ higher decoding throughput over Qwen2-VL.
@@ 4.1 Temporal Localization
Table 10: Temporal localization. LITA results are from their original paper, while VILA-1.5 results are based on our reproduction. Our NVILA uses the same data mixture as VILA-1.5; the only difference is the backbone VLM.
ActivityNet-RTL
#Frames Mean IoU Precision@0.5
LITA 7B 100 24.1 21.1
LITA 13B 100 28.6 25.9
VILA-1.5 8B 256 32.1 29.3
NVILA 8B 256 34.8 32.1
Figure 6: Robotic navigation. NVILA deployed as a Vision-Language Navigation agent, navigating environments using language instructions and visual observations (Top: simulation, Bottom: real-world). The real-world setup features a Unitree Go2 robot equipped with a LiDAR sensor at the base of its head and an Intel RealSense Camera mounted on top. On the server side, an RTX 4090 GPU powers the NVILA-8B model, configured with an 8-frame context length for action generation.
@@ 4.2 Robotic Navigation
Table 11: Robotic navigation. All numbers are from NaVILA, except for those of NVILA. All models are provided with only RGB inputs. We refer the readers to NaVILA [8] for more details.
@@
Obs. NE $\downarrow$ OS $\uparrow$ SR $\uparrow$ SPL $\uparrow$
Seq2Seq RGB 10.10 8.0 0.0 0.0
CMA RGB 9.55 10.0 5.0 4.0
NaVid 7B RGB 5.47 49.0 37.0 35.0
NVILA 8B RGB 5.43 60.4 53.3 48.8
4.3 Medical Application
NVILA also offers transformative potential in the medical domain. Such integration promises advancements in diagnostic accuracy, clinical decision-making, and data interpretation. The NVILA-M3 framework [11] introduces a novel approach by integrating multiple domain-expert models tailored to specific medical tasks, such as image segmentation and classification. These expert models are designed to extract and interpret intricate features that are otherwise difficult for general VLM’s to discern. By coupling these specialized models with a vision-language learning paradigm, NVILA-M3 achieves enhanced performance, facilitating the learning of nuanced relationships between visual inputs and their textual annotations. This integration not only improves task-specific outcomes but also sets a foundation for the development of more robust and context-aware VLMs in the healthcare domain. NVILA-M3 indicated that an overall improvement of 9% can be achieved via usage of expert models over existing SOTA, a few key results can be observed in Table. 12. This underscores the importance of leveraging domain expertise to bridge the gap between generalized AI capabilities and the demands of specialized applications, demonstrating the potential for VLMs to revolutionize fields where precision and specificity are paramount.
Table 12: Medical application. Performance of best M3 model on key benchmarks is shown. Task-specific SOTA baselines and datasets are described in the experiments section [11]. Metrics for VQA is accuracy, for report generation BLEU-4 & ROUGE and for classification F1 score have been utilized
VQA Report Gen. Classif.
Rad Path CXR CheXpert
Med-Gemini 78.8 83.3 20.5 28.3 48.3
VILA-M3 8B 84.7 91.0 21.1 32.0 61.6
NVILA 8B 85.5 92.9 22.8 32.8 61.1
Task-spfc. SOTA 84.2 91.7 15.4 30.6 51.5
@@ 5.1 Visual Language Models
Meanwhile, open-source VLMs continue to evolve, improving at both the system/framework level [73] and the algorithm/recipe level [2], progressively narrowing the performance gap between proprietary and open-source models [19, 74, 75, 76, 5]. These recent advancements have led many open VLM models to claim performance levels comparable to, or even exceeding, leading proprietary models such as GPT-4V and GPT-4o. Some representative examples include InternVL2 [3], Qwen2-VL [5], LLaVA-OneVision [4], Llama 3.2 Vision [77], Molmo [30], NVLM [76], and MiniCPM-V [18].